8. Convert into mole :
(a) 12 g of oxygen gas
(b) 20 g of water
(c) 22 g of carbon dioxide

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(a) 32 g of oxygen gas = 1 mole Then, 12 g of oxygen gas = 12/32 mole = 0.375 mole (b) 18 g of water = 1 mole, then, 20 g of water = 20/18 mole = 1.111 mole (c) 44 g of carbon dioxide = 1 mole, then, 22 g of carbon dioxide = 22/44 mole = 0.5 mole